Consider a Multi-Index Model (MIM) specification for the portfolio return:


π‘Ÿπ‘π‘‘ = 𝛼𝑝+𝛽𝑝1𝐹1𝑑 + 𝛽𝑝2𝐹2𝑑 + πœ€π‘π‘‘


a) Derive the functional form for the variance of π‘Ÿπ‘π‘‘, denoted as πœŽπ‘2.


b) In deriving πœŽπ‘2, what are the key assumptions you made under the MIM?


c) If you estimate the above MIM as a regression, and you find that the variance of the residual return πœ€π‘π‘‘ represents a substantial portion of πœŽπ‘2 i.e. low 𝑅2 , how

would you interpret this finding? [Hint: More than 1 reason.]


1
Expert's answer
2021-04-27T07:12:40-0400

a)

"r_{p𝑡}=\alpha _{p𝑡}+\beta_{p1}F_{1𝑡}+\beta _{p2}F_{2t}+\xi _{p𝑡}"


To remove relation between F1 and F2, the coefficient of the following equation can be derived by regression analysis.

"F_2=e_o+e_1F_1+d_i"

where: eo and e1 = the regression coefficients

di = the random error term

By the assumptions of regression analysis, is uncorrelated with. Therefore:

"F_2=e_o+e_1F_1+d_i"

Which is an index of performance of the sector index without the effect of F1

defining:

"\u02c6d_i=F_2-(\u02c6e_o+\u02c6e_1F1)"


an index is obtained that is uncorrelated with the market. By solving for F2 and substituting

"r_{p𝑡}=\alpha _{p𝑡}+\beta_{p1}F_{1𝑡}+\beta _{p2}F_{2t}-\beta_{p2}ˆe_o-\beta _{p2}ˆe_1F_{1t}+\xi_{pt}"



"r_{p𝑡}=(\alpha_{pt}-\beta_{p2}ˆe_o)+(\beta_{p1}-\beta_{p2}ˆe_1)F_1+\beta_{p2}F_2+\xi_{pt}"


b)

By assumption

"E[(F_{1t}+\u02c6F)(F_{2t}+\u02c6F_2)]=0"

"E[(F_{1t}+\u02c6F_1)e_i]=0"

"E[(F_{2t}+\u02c6F_2)e_i]=0"

and


"E(F_{1t}+ˆF_1)^2=\sigma _1^2"

"E(F_{2t}+ˆF_2)^2=\sigma_2^2"

"E(e_1^2)=\sigma _{ei}^2"


therefore

"\sigma_{pt}^2=\beta_{p1}^2\sigma_1^2+b_{p2}\sigma_2^2+\sigma_{ep}^2"


c)

covariance between security pt and j can be expressed as:

let "pt=i"

"\beta_p=b_i"

"F=I"

"\xi_{pt}=C_i"





substituting the answer becomes

"\beta_{pt1}\beta_{j1}\sigma_1^2+\beta_{pt2}\beta_{j2}\sigma_2^2"